# Are They the Same? Exploring Visual Correspondence Shortcomings of Multimodal LLMs

### v2

@@ Abstract
Recent advancements in multimodal models have shown a strong ability in visual perception, reasoning abilities, and vision-language understanding. However, studies on visual matching ability are missing, where finding the visual correspondence of objects is essential in vision research. Our research reveals that the matching capabilities in recent multimodal LLMs (MLLMs) still exhibit systematic shortcomings, even with current strong MLLMs models, GPT-4o. In particular, we construct a Multimodal Visual Matching (MMVM) benchmark to fairly benchmark over 30 different MLLMs. The MMVM benchmark is built from 15 open-source datasets and Internet videos with manual annotation. We categorize the data samples of MMVM benchmark into eight aspects based on the required cues and capabilities to more comprehensively evaluate and analyze current MLLMs. In addition, we have designed an automatic annotation pipeline to generate the MMVM SFT dataset, including 220K visual matching data with reasoning annotation. Finally, we present CoLVA, a novel contrastive MLLM with two novel technical designs: fine-grained vision expert with object-level contrastive learning and instruction augmentation strategy. CoLVA achieves 51.06% overall accuracy (OA) on the MMVM benchmark, surpassing GPT-4o and baseline by 8.41% and 23.58% OA, respectively. The results show the effectiveness of our MMVM SFT dataset and our novel technical designs. Code, benchmark, dataset, and models are available at https://github.com/zhouyiks/CoLVA.
Figure 1: Visualization results of GPT-4o and our proposed CoLVA on challenging cases of MMVM benchmarks. GPT-4o’s answers are incorrect for all these examples, with the errors highlighted in red. The correct answers in the options are highlighted in green.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) [48, 50, 12, 81, 58] have made remarkable progress with the development of LLMs [76, 29, 91]. They have greatly benefited various applications, including image and video understanding [107, 48, 6], visual question answering (VQA) [12, 81], and visual grounding [31, 65, 107].
@@
Despite the advancements of MLLMs with various capabilities [6, 35, 12, 81, 41, 69, 24], they often struggle with visual correspondence, a fundamental ability that plays a key role in several vision tasks, including tracking [66, 36], feature matching [5], and reconstruction [21]. As shown in Fig. 1, even the GPT-4o [58] cannot understand some simple matching questions well. This limitation is critical, as it hinders MLLMs from comprehending correspondence-aware information.
@@
Based on this motivation, we aim to systematically analyze this problem in MLLMs [48, 58, 12, 81] and propose a corresponding method to address it. At first, building a new and challenging benchmark on instance-level correspondence across multiple images is required due to the lack of comprehensive evaluations for this direction. Specifically, we collect 1,510 samples from 15 public datasets [61, 92, 80, 16, 97, 3, 14, 27, 25, 60, 57, 17, 56, 7, 78] and internet video platforms. These samples encompass various scenes, such as indoor environments, urban settings, and card and billiards games. Each sample is carefully annotated with multi-image QA pairs by three skilled annotators. For the benchmark, we evaluate 36 state-of-the-art (SOTA) MLLM methods across eight aspects, including color, shape, posture, size, textual or logo markers, and more. The quantitative evaluation on our benchmark highlights the merits of our work, as the best public model, Qwen2-VL-72B-Instruct [81], achieves only 38% accuracy. Additionally, as shown in Fig. 2, based on evaluation results across different aspects, open-source MLLMs [12, 53, 34, 81] perform significantly worse on samples requiring complex cues for reasoning (such as relative position (RP), object orientation (OO), binding relationship with other objects (BR), and object markers (OM)) compared to samples that can be completed using only appearance cues (such as size (SZ), textual or LOGO markers (TM), shape or posture (SP), and color (CL)).
Despite the unsatisfactory results achieved by state-of-the-art (SOTA) MLLM models [53, 52, 12, 81, 58, 34], we recognize the potential of existing architectural designs for correspondence-aware learning by analyzing the attention map between queries and candidate objects. This insight motivates us to develop an automatic data generation pipeline for building a high-quality visual matching SFT dataset, named MMVM. We propose CoLVA with two techniques to help with correspondence training: fine-grained vision expert with object-level contrastive learning and instruction augmentation. On the one hand, the MMVM dataset includes multi-choice QA pairs, each accompanied by textual reasoning and explanation. On the other hand, we construct a strong baseline CoLVA by integrating two simple but effective techniques into the InternVL2 [12] and fine-tuning it with our MMVM dataset. These tricks enhance the model’s instance-discriminative representation. Finally, extensive experiments demonstrate the effectiveness of our MMVM dataset and network design, achieving 32.25% and 7.22% improvements over the baseline and the closed GPT-4o. To sum up, our contributions are three-fold:
  • We establish a challenging benchmark for the correspondence problem in Multimodal LLMs, which has been insufficiently explored. Our analysis reveals that current MLLM methods are inadequate for addressing this issue.
  • We propose a high-quality MMVM dataset and two simple, yet effective techniques for correspondence learning designed to enhance the instance-level discriminative ability.
@@
  • Extensive experiments demonstrate the effectiveness of our proposed dataset. Moreover, our proposed techniques significantly improve the performance of existing MLLMs.
Figure 2: The SOTA open-source MLLMs’ performance on the MMVM benchmark. Our benchmark evaluates visual matching tasks from 8 aspects. SZ, TM, SP, and CL samples can be completed using appearance cues, while RP, OO, BR, and OM samples require complex cues and reasoning. Our CoLVA surpasses current SOTA MLLMs in all aspects of the MMVM benchmark.
@@ 2 Related Work
Figure 3: Visualization of MMVM Benchmark. Our MMVM Benchmark contains 1,510 manually annotated multi-image QA pairs, 8 matching patterns, and 2 types of object referring methods. We collect those evaluate samples from 15 open-source video datasets and various internet video platforms.
Visual Corresponding Learning. Learning instance discriminate features is critical to many applications, including object tracking, person re-identification, and multi-view reconstruction. Several works [111, 55, 50, 81] explore the cross-image understanding of MLLMs, and most works follow the VQA pipeline. Our works are inspired by previous visual corresponding learning [85, 104, 105, 36, 37, 110, 106, 103] and present a new learning framework with contrastive visual tokens for current MLLMs.
@@
Figure 5: The attention map of image pairs when the MLLMs generate responses. In all image pairs, a query object is specified in the first image, while multiple candidate objects are specified in the second image. The MLLMs select the candidate object in the second image that matches the query object.
3 MMVM Benchmark and Dataset
@@ 3.1 MMVM Benchmark
Dataset Format. To ensure the absolute fairness of the benchmark, instead of using ChatGPT as a judge, we formulate instruction examples into multiple-choice questions, similar to MMBench [51]. We use text prompts or visual prompts to specify objects. Considering that most MLLMs cannot understand additional visual prompts, we overlay the visual prompts onto the images using highlight contours of different colors and a number tag. The evaluation item consists of image pairs (more than two images), a question, and options. The MLLM must select the correct answer from the given options based on the question and image pairs.
Data Collection. Our benchmark samples contain two parts: one is filtered from public datasets, and the other is manually selected from internet videos. Firstly, we carefully filter the most challenging 790 samples from 15 video segmentation, tracking, and multi-view matching datasets (including OVIS [61], Youtube-VIS 2021 [92], LVVIS [80], MOSE [16], BDD100k [97], BURST [3], SportMOT [14], VideoCube [27], MultiviewX [25], VastTrack [60], MOT17 [57], LaSOT [17], VIPSeg [56], WildTrack [7], and OxUvA [78]), and manually annotate the QA pairs. These samples include various scenes such as indoor, urban, and wilderness environments and contain images captured by different platforms, including fixed cameras, mobile phones, vehicle-mounted cameras, and drones. Secondly, to further enhance the diversity and challenge of the MMVM benchmark, we manually collect 720 videos from various internet video platforms. These videos include more challenging scenarios such as crowded streets, card games, and billiards games. For these samples, we also carefully annotate QA pairs manually. For some samples, we use text prompts rather than visual prompts to refer to objects to more comprehensively benchmark MLLMs’ visual matching performance. Combining the above two sample parts, we construct the MMVM benchmark, which includes 1,510 manually annotated multi-image QA pairs. Finally, these 1,510 samples are cross-validated by three skilled annotators to ensure complete correct. For more annotated examples, please see our appendix.
@@
Evaluation Metric. We use accuracy as the evaluation metric, following previous VQA benchamrks [18, 22]. To analyze the deficiencies of various MLLMs in visual matching more thoroughly, we manually categorized all samples based on the cues for matching.
The right side of Fig. 3 shows samples of different categories: 1) Color (CL). 2) Shape or posture (SP). 3) Textual or LOGO markers (TM). 4) Size (SZ). 5) Relative position in the scene (RP). 6) Object orientation and movement (OO). 7) Binding relationship with other objects (BR). 8) Object Markers (OM).
The samples of CL, SP, TM, and SZ rely on MLLMs’ ability to perceive basic object information such as color, size, position, posture, and OCR capabilities. The samples of RP, OO, BR, and OM require MLLMs to reason for vision matching through surrounding environments or objects.
3.2 MMVM Dataset
To improve the visual matching performance of current MLLMs, we collect the MMVM dataset, a large-scale visual matching dataset.
Multiple-choice QA Generation. We filter and reorganize the train sets of current video segmentation datasets, including OVIS [61], YouTube-VIS 2021 [92], VIPSeg [56], BDD100K [97], and BURST [3]. As shown on the left side of Fig. 4, we sample frames at fixed 1-second intervals for each video, then organize adjacent frames into image pairs. Each object pair will generate a question-answer (QA) pair, with incorrect options being other objects in the images. Following this process, we have generated 220K QA pairs.
Reason Annotations. Multiple-choice training data can hardly provide strong supervision for MLLMs. Inspired by chain-of-thought [83], we append reasoning and explanation for each multiple-choice question. For this purpose, we design a pipeline to prompt MLLMs to generate reasons automatically. As shown in the middle of Fig. 4, first, we prompt the SOTA MLLM InternVL2-76B [12] to annotate basic information for all query and candidate objects, including color, size, position, posture, etc. Then, we give both the answer (which two objects are the same) and the objects’ basic information as conditions and prompt InternVL2-76B to generate corresponding reasons. Finally, we obtain 220K matching QA pairs with reasons, with one data sample shown on the right side of Fig. 4.
@@ 4.1 Analysis of Current MLLMs’ Shortcomings
We have evaluated 36 SOTA MLLMs on the MMVM benchmark. However, the results are promising: None of the MLLMs can achieve an overall accuracy exceeding 50% (Tab. 1). This phenomenon suggests significant deficiencies in current MLLMs’ performance on the visual matching task. We argue that two main factors cause this phenomenon: 1) Although current MLLMs possess a certain capability to recognize objects’ appearances and positions, they lack the corresponding data to teach them how to utilize this basic knowledge and these abilities for visual matching, even in a simple sense. This hypothesis stems from two observations: Firstly, as shown in Fig. 5, we visualize the attention maps between multiple SOTA MLLMs’ responses and images, which show that MLLMs correctly focus on the query objects. Secondly, we find that these MLLMs, such as InternVL2-76B, can accurately identify query objects’ basic information, including category, color, position, shape, posture, and relationships with other objects in the environment. However, it only achieved a score of 25 on the MMVM benchmark. These results indicate that current MLLMs possess the necessary capabilities for visual matching, yet they cannot properly utilize these abilities to perform visual matching successfully. 2) Current MLLMs rely on CLIP models to understand images and cannot comprehend fine-grained and discriminative visual features, which are essential for visual matching since candidate objects often share extremely similar semantic information. As shown in Fig. 5, the attention maps demonstrate that MLLMs give almost equal attention to all candidate objects, unable to distinguish the correct object.
@@ 4.2 CoLVA
To address the shortcomings summarized in Sec. 4.1, we propose a novel Object-level Contrastive Learning (OCL) strategy and introduce a fine-grained vision expert to provide the discriminative and fine-grained visual features, thereby improving the MLLM’s visual matching performance. Additionally, we propose an instruction augmentation strategy to facilitate MLLM training. These two novel designs will be detailed in the following.
@@
Object-Level Contrastive Learning. Inspired by the success of contrastive learning in visual representation [23, 10, 59], tracking [89, 90], and video segmentation [85, 105, 37, 95, 36], we introduce a novel object-level contrastive learning (OCL) strategy to help MLLM learn more discriminative features for better visual matching.
@@
This instruction format allows the MMVM data to be seamlessly compatible with InternVL2 [12] for direct training. However, it still has some drawbacks. 1) Editing the image may disrupt the original object information, especially for small objects. 2) Since the tags are used indirectly to refer to objects, gradients cannot be directly backpropagated to the corresponding image features. To address these challenges, we designed a new instruction format to support direct use of object-level representations to refer to objects:
@@
where <Object_Info> is: “object-1: <Obj 1>, object-2: <Obj 2>, …, object-n: <Obj n>", with “<Obj 1>" to “<Obj n>" replaced by the respective object-level representations. This instruction format allows gradients to be directly backpropagated through the object-level representations to the corresponding image features, enabling the MLLM to learn the required discriminative and fine-grained features more effectively.
During the training, we randomly use these two instruction formats to organize the data, which we refer to as instruction augmentation. Only the first instruction format that editing the images is adopted to support fair comparisons across all MLLMs.

### v3

@@ Abstract
Recent advancements in multimodal large language models (MLLM) have shown a strong ability in visual perception, reasoning abilities, and vision-language understanding. However, the visual matching ability of MLLMs is rarely studied, despite finding the visual correspondence of objects is essential in computer vision. Our research reveals that the matching capabilities in recent MLLMs still exhibit systematic shortcomings, even with current strong MLLMs models, GPT-4o. In particular, we construct a Multimodal Visual Matching (MMVM) benchmark to fairly benchmark over 30 different MLLMs. The MMVM benchmark is built from 15 open-source datasets and Internet videos with manual annotation. In addition, we have designed an automatic annotation pipeline to generate the MMVM SFT dataset, including 220K visual matching data with reasoning annotation. To our knowledge, this is the first visual corresponding dataset and benchmark for the MLLM community. Finally, we present CoLVA, a novel contrastive MLLM with two novel technical designs: fine-grained vision expert with object-level contrastive learning and instruction augmentation strategy. The former learns instance discriminative tokens, while the latter further improves instruction following ability. CoLVA-InternVL2-4B achieves an overall accuracy (OA) of 49.80% on the MMVM benchmark, surpassing GPT-4o and the best open-source MLLM, Qwen2VL-72B, by 7.15% and 11.72% OA, respectively. These results demonstrate the effectiveness of our MMVM SFT dataset and our novel technical designs. Code, benchmark, dataset, and models will be released.
Figure 1: Visualization results of GPT-4o and our proposed CoLVA on challenging cases of MMVM benchmarks. The GPT-4o’s answers are incorrect for all these examples, with the errors highlighted in red. The correct answers in the options are highlighted in green.
00footnotetext: *Equal contribution. Corresponding author. Project leader.
@@ 1 Introduction
MLLMs [50, 52, 12, 83, 60] have made remarkable progress with the development of Large Language Models (LLMs) [78, 30, 93]. They have greatly benefited various applications, including image and video understanding [109, 50, 6], visual question answering (VQA) [12, 83], and visual grounding [32, 67, 109]. Despite the advancements of MLLMs with various capabilities [6, 36, 12, 83, 43, 71, 25], they often struggle with visual correspondence, a fundamental ability that plays a key role in several vision tasks, including tracking [68, 37], feature matching [5], and reconstruction [22]. As shown in Fig. 1, even the GPT-4o [60] cannot understand some simple matching questions well. This limitation is critical, as it hinders MLLMs from comprehending correspondence-aware information.
@@
Based on this motivation, we aim to systematically analyze this problem in MLLMs [50, 60, 12, 83] and propose a corresponding method to address it. First, a new and challenging benchmark on instance-level correspondence across multiple images is required due to the lack of comprehensive evaluations for this direction. Specifically, we collect 1,510 samples from both 15 public video datasets [63, 94, 82, 16, 99, 3, 14, 28, 26, 62, 59, 17, 58, 7, 80] and internet video platforms. These samples encompass various scenes, including indoor environments, urban settings, cartoons, drone footage, and various social activity scenarios. Each sample is meticulously annotated with multi-image QA pairs by three skilled annotators. The diversity of these samples enables us to evaluate the visual matching capabilities of MLLMs across multiple dimensions of matching cues. We summarize eight types of matching cues (such as color, markers), which are the most frequently encountered by humans. (See the Sec. 3)
Then, we evaluate 36 state-of-the-art (SOTA) MLLM methods on our benchmark. The quantitative evaluation in our benchmark highlights the merits of our work, as the strong model, Qwen2-VL-72B-Instruct [83], achieves only 38% overall accuracy. This indicates that current state-of-the-art (SOTA) MLLMs exhibit notable matching shortcomings. Through quantitative experiments and PCA visualization analysis (Fig. 2), we identify two primary factors contributing to these visual shortcomings: 1) Although current MLLMs possess a specific capability to recognize objects’ appearances and positions, they lack the corresponding data to teach them how to utilize this foundational knowledge and these abilities for visual matching; 2) Current MLLMs rely on CLIP models to understand images and cannot comprehend fine-grained and discriminative visual features.
Figure 2: The PCA visualization of learned object embeddings by InternVL2-4B and our CoLVA-4B. The object embeddings are obtained by applying average pooling to the visual tokens using mask annotations. The red star represents the query object in the first image. The red dot represents the matched target in the second image. The blues dots represent other candidates. More PCA visualizations can be found in the appendix.
These findings motivate us to develop an automatic data generation pipeline for building a high-quality visual matching SFT dataset (MMVM dataset). The MMVM dataset includes 220k multi-choice QA pairs. Each is accompanied by matching rationales. We establish a simple yet effective baseline, CoLVA, and fine-tune it using our MMVM dataset. CoLVA integrates two simple yet effective techniques into existing SOTA MLLMs, such as InternVL2 [12] and Qwen2VL [83], to enhance correspondence training: a fine-grained vision expert with object-level contrastive learning (OCL) and instruction augmentation (IA). Specifically, we perform object-level contrastive learning between the MLLM visual encoder and the vision expert, motivated by previous works [18, 38]. This enables the vision expert to learn discriminative features within the semantic space of the MLLM. First, it preserves fine-grained visual features since our benchmark involves detailed visual prompts as inputs. In addition, it achieves modality alignment through contrastive learning. This dual-purpose design highlights the novelty of our OCL strategy. Furthermore, we integrate the learned discriminative object-level features into the instructions. This allows gradients to be directly backpropagated through the object-level features to the corresponding image features, enabling the MLLM to learn the required discriminative and fine-grained features more effectively. Moreover, this enhances our ability to refer to multiple objects within the images. Finally, extensive experiments demonstrate the effectiveness of our MMVM dataset and network design. Our CoLVA-InternVL2-4B achieves improvements of 11.72% and 7.15% over the open-source Qwen2VL-72B and the proprietary GPT-4o, respectively.
To sum up, our contributions are four-fold:
  • We establish a challenging benchmark for the visual matching problem in Multimodal LLMs.
  • We propose a high-quality MMVM dataset, which contains 220k matching QA pairs with reasoning texts.
  • We propose two simple, yet effective techniques for correspondence learning.
@@
  • Extensive experiments demonstrate the effectiveness of our proposed dataset and technical contributions.
@@ 2 Related Work
Figure 4: The statistics of the MMVM benchmark. The left side presents the statistics of the example counts for matching cues, while the right side displays the statistics regarding the number of candidates in the examples.
Figure 5: Visualization of MMVM Benchmark. Our MMVM Benchmark contains 1,510 manually annotated multi-image QA pairs, 8 matching patterns, and 2 types of object referring methods. We collect the evaluation samples from 15 open-source video datasets and various internet video platforms.
Visual Corresponding Learning. Learning instance discriminative features is critical to many applications, including object tracking, person re-identification, and multi-view reconstruction. Several works [113, 57, 52, 83] explore the cross-image understanding of MLLMs, and most works follow the VQA pipeline. Our works are inspired by previous visual corresponding learning [87, 106, 107, 37, 39, 112, 108, 105] and present a new learning framework with contrastive visual tokens for current MLLMs.
@@
3 MMVM Dataset and Benchmark
We first introduce the strategy for constructing the MMVM dataset (detailed in Sec. 3.1). We then detail the MMVM benchmark in Sec. 3.2.
3.1 MMVM Dataset
To construct a large-scale visual matching dataset, we leverage existing video datasets to generate multiple-choice QA pairs (Step 1) and collect reasoning texts for visual matching by prompting advanced MLLMs (Step 2). Finally, we organize the multi-choice QA pairs and reasoning texts into a multi-turn dialogue format (Step 3).
Multiple-choice QA Generation. We filter and reorganize the train sets of current video segmentation datasets, including OVIS [63], YouTube-VIS 2021 [94], VIPSeg [58], BDD100K [99], and BURST [3]. As illustrated on the left side of Fig. 3, we sample frames at fixed 1-second intervals for each video and subsequently organize adjacent frames into image pairs. Each image pair contains multiple objects, enabling the generation of multiple visual matching QA pairs. We directly utilize existing mask annotations to refer to objects and form multiple candidate options, while leveraging existing matching annotations to construct the answers. For image pairs lacking mask annotations, we employ SAM [31] to automatically generate the corresponding mask annotations. In cases where image pairs lack matching annotations, we utilize the Re-identification method [37] to obtain the matching relationships between objects. Thus, we generate 220K QA pairs in total, the entire process is shown in left side of Fig. 3.
Reason Generation. Multiple-choice training data can hardly provide text supervision for MLLMs. Inspired by chain-of-thought [85], we append reasoning and explanation for each multiple-choice question. For this purpose, we design a pipeline to prompt MLLMs with mask annotations and matching annotations to generate reasons automatically. Although our experiments indicate that existing MLLMs exhibit poor visual matching capabilities, within our pipeline, MLLMs are not required to perform visual matching themselves. Instead, they only need to summarize visual cues that are beneficial for visual matching based on the provided annotations. As shown in the middle of Fig. 3, first, we prompt the stronger MLLM InternVL2-76B [12] to annotate basic information for all query and candidate objects, including color, size, position, posture, etc. Then, we give both the answer (which two objects are the same) and the objects’ basic information as conditions and prompt InternVL2-76B to generate corresponding reasons. Finally, we obtain 220K matching QA pairs with reasons.
Match VQA Formulation. As illustrated on the right side of Fig. 3, we organize the multiple-choice QA pairs and reasoning texts generated in Step 1 and Step 2 into a two-turn dialogue. The first turn requires the model to make a selection, whereas the second turn requires the model to provide a reasoning for its chosen answer. Ultimately, we obtain a dataset comprising 220k multi-turn dialogue samples.
@@ 3.2 MMVM Benchmark
To evaluate the visual matching capabilities of MLLMs, we also collect image pairs from internet video platforms and the validation splits of existing video datasets. These pairs are manually annotated with mask annotations and matching annotations by three experts. We specifically select challenging examples to form our MMVM benchmark, which comprise a total of 1,510 examples.
Example Format. As shown in Fig. 5, we use text prompts or visual prompts to specify objects. Considering that most MLLMs cannot understand additional visual prompts, we overlay the visual prompts onto the images using highlight contours of different colors and a number tag. Each example consists of image pairs (more than two images), a question, and options. The MLLM must select the correct answer from the given options based on the question and image pairs.
Benchmark Statistics. Our MMVM benchmark comprise a total of 1,510 examples. Among these, 790 examples are derived from 15 video segmentation, tracking, and multi-view matching datasets (including OVIS [63], Youtube-VIS 2021 [94], LVVIS [82], MOSE [16], BDD100k [99], BURST [3], SportMOT [14], VideoCube [28], MultiviewX [26], VastTrack [62], MOT17 [59], LaSOT [17], VIPSeg [58], WildTrack [7], and OxUvA [80]). To mitigate potential overlap with MMVM training data (Sec. 3.1), we exclusively selected image pairs from the validation splits of these datasets and re-annotated them manually. To further augment the diversity and complexity of the MMVM benchmark, we manually gathered 720 videos from a variety of internet video platforms. The diversity of these 1,510 examples enables us to evaluate the ability of MLLMs to discern and comprehend multiple matching cues. As illustrated in Fig. 4, we enumerate eight types of matching cues, including: 1) Color (CL), 2) Shape or posture (SP), 3) Textual or LOGO markers (TM), 4) Size (SZ), 5) Relative position in the scene (RP), 6) Object orientation and movement (OO), 7) Binding relationship with other objects (BR), and 8) Object Markers (OM). The examples of these matching cues are shown in Fig. 5. Each example may exhibit multiple matching cues, but we annotate only the most salient ones. CL and RP are the most prevalent cues. Each example includes multiple candidate options, as depicted in Fig. 4. The minimum, maximum, and average number of choices per example are 2, 37, and 10, respectively.
@@
Evaluation Metric. Following previous works [19, 23], we adopt accuracy as the main evaluation metric. In addition to calculating an overall accuracy, we also separately assess accuracy for eight distinct matching cues.
@@ 4.1 Analysis of Current MLLMs’ Shortcomings
We have evaluated multiple SOTA open-source and proprietary MLLMs on the MMVM benchmark. However, the results reveal a notable observation: none of the open-source MLLMs achieve an overall accuracy exceeding 50% (Tab. 1). This phenomenon suggests significant deficiencies in current MLLMs’ performance on the visual matching task. We argue that two main factors cause this phenomenon: 1) Although current MLLMs possess a certain capability to recognize objects’ appearances and positions, they lack the corresponding data to teach them how to utilize this basic knowledge and these abilities for visual matching, even in a simple sense. This hypothesis is supported by two observations: First, in the annotation pipeline of the MMVM dataset, InternVL2 can accurately identify the basic information of query objects, yet it achieves a notably low score on the MMVM benchmark, as illustrated in Tab. 1. Second, when we fine-tune InternVL2 using our MMVM dataset enriched with matching reasoning, its performance improves significantly (+14.76%). 2) Current MLLMs rely on CLIP models to understand images and cannot comprehend fine-grained and discriminative visual features, which are essential for visual matching since candidate objects often share extremely similar semantic information. As illustrated in Fig. 2, we conduct a PCA visualization analysis on the object embeddings learned by InternVL2. The results show that the matched target (represented by a red dot) and other candidate objects (represented by blue dots) are clustered together, while being distant from the query object (represented by a red star). This clustering pattern makes it challenging for MLLMs to distinguish the correct object.
@@ 4.2 CoLVA
To address the shortcomings summarized in Sec. 4.1, we propose a novel Object-level Contrastive Learning (OCL) strategy and introduce a fine-grained vision encoder to provide the discriminative and fine-grained visual features, thereby improving the MLLM’s visual matching performance. Additionally, we propose an instruction augmentation strategy to facilitate MLLM training. These two novel technical designs will be detailed in the following.
@@
Object-Level Contrastive Learning. Inspired by the success of contrastive learning in visual tracking [24, 10, 61], tracking [91, 92], and video segmentation [87, 107, 39, 97, 37], we introduce a novel object-level contrastive learning (OCL) strategy to help MLLM learn more discriminative features for better visual matching. Unlike previous contrastive learning approaches that learning features using shared tracking heads, our method conducts contrastive learning between two distinct vision encoders: the MLLM visual encoder and an additional visual expert, as illustrated on the left side of Fig. 6. This design allows the visual expert to learn discriminative features within the semantic space of the MLLM. On the one hand, it preserves fine-grained visual features, while on the other hand, it achieves modality alignment through contrastive learning. We employ the OCL strategy during the pre-training phase. As shown in Tab. 5, the OCL strategy outperforms other pre-training methods.
@@
This instruction format allows the MMVM data to be seamlessly compatible with InternVL2 [12] for direct training. However, it still has some drawbacks. 1) Editing the image may disrupt the original object information, especially for small objects. 2) Since the tags are used indirectly to refer to objects, gradients cannot be directly backpropagated to the corresponding image features. To address these problems, we designed a new instruction format to support direct use of object-level representations to refer to objects:
@@
where <Object_Info> is: “object-1: <Obj 1>, object-2: <Obj 2>, …, object-n: <Obj n>", with “<Obj 1>" to “<Obj n>" replaced by the respective object-level representations. This instruction format allows gradients to be directly backpropagated through the object-level representations to the corresponding image features, enabling the MLLM to learn the required discriminative and fine-grained features more effectively. We randomly use these two instruction formats to organize the data, which we refer to as instruction augmentation.